OpenStudy (anonymous):

solve and check.... 2n - 7 = 15

OpenStudy (albert0898):

2n - 7 = 15 Add seven to both sides 2n = 22 Divide both sides by 2 n = 11

OpenStudy (kropot72):

Add 7 to both sides: 2n - 7 + 7 = 15 +7 2n = 22 n = 11